Solve the two equations below by elimination
3x + 5y = 7
x + y = 2
please help!!

Answer by mananth(16949)

3 x + 5 y = 7 .............1
1 x + 1 y = 2 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -5
3 x + 5 y = 7
-5 x + -5 y = -10
Add the two equations
-2x=-3
/-2
x=1.50
plug value of x in (1)
3x+5y=7
4.5+5y=7
5 y = 7-4.5
5 y = 2.5
y = 0.50